# Reprogramming Pretrained Target-Specific Diffusion Models for Dual-Target Drug Design

### v1

@@ 1 Introduction
A promising paradigm of rational drug design is structure-based drug design (SBDD) [1], which uses computational chemistry tools in which the 3D structure of a protein target is used as the basis to identify or design new chemical entities. The foundation of structure-based drug design has been grounded in the lock-and-key hypothesis, positing that an optimal ligand molecule should possess a structure that is complementary to the target site. Recently, dual-target drug design, which aims to design “one key” for “two locks”, has attracted significant attention. Precisely, dual-target drug design [4] is a strategy in pharmaceutical research that aims to design single ligand molecules capable of interacting with two different biological targets simultaneously. A dual-target drug can potentially lower the odds of resistance developing [51] and effectively manage the disease which involve complex biological pathways with multiple proteins [40]. Recent years have witnessed a noticeable increase in the FDA’s approval of dual-target drugs [27, 29].
@@ 2 Related Work
Molecular linker design, which enables the connection of molecular fragments to form potent compounds, is an effective approach in rational drug discovery. Approaches like DeLinker [21] and Develop [22] design linkers by utilizing molecular graphs with distance and angle information between anchor atoms, but they lack 3D structural information of molecules. More recent techniques, such as 3DLinker [17] and DiffLinker [20], generate linkers directly in 3D space using conditional VAEs and diffusion models, respectively, but they assume known fragment poses. LinkerNet [12] relaxes this assumption by co-designing molecular fragment poses and the linker, making it applicable in cases where fragment poses are unknown, such as in the linker design of PROTACs (PROteolysis TArgeting Chimeras). Since pharmacophore combination is a traditional strategy to design dual-target drugs, we repurpose linker design methods as strong baselines for dual-target drug design.
@@ 3.1 Data Curation
To collect drug combination pairs, we start from DrugCombDB [^1] http://drugcombdb.denglab.org/main [30]. DrugCombDB is a comprehensive database devoted to the curation of drug combinations from various data sources including high-throughput screening (HTS) assays, manual curations from the literature, FDA Orange Book and external databases. DrugCombDB comprises a total of 448,555 combinations of drugs, encompassing 2,887 unique drugs and 124 human cancer cell lines. Particularly, DrugCombDB has more than 6,000,000 quantitative dose responses, from which we determine whether a drug combination is synergistic or not. Specifically, a drug combination with positive zero interaction potency (ZIP), Bliss, Loewe and the highest single agent (HSA) scores simultaneously in at least one cell line is supposed to be a synergistic one. Please refer to Appendix A for a comprehensive understanding of these scores.
@@ 3.2 Reprogramming Target-Specific Diffusion Models for Dual-Target Drug Design
We name the compositional reverse sampling with the above SE(3)-equivariant message at each layer as DualDiff. The proof of SE(3)-equivariance can be found in Appendix D. This more meticulous composition is supposed to lead to higher-quality samples.
@@ 3.3 Repurposing Linker Design Methods for Dual-Target Drug Design
Pharmacophore combination is a prevalent strategy in traditional dual-target drug design, requiring the specialized knowledge of chemists. To automate this procedure, We design a strategy to identity crucial fragments from reference molecules of dual targets in our dataset and apply linker design methods [20, 12] to link the fragments and obtain complete molecules.
@@ 4.1 Experimental Setup
We compare method with various baselines: Pocket2Mol [37] generates 3D molecules atom by atom in an autoregressive manner given a specific protein binding site. TargetDiff [14] is a diffusion-based method which generates atom coordinates and atom types in a non-autoregressive way. Note that Pocket2Mol and TargetDiff are both proposed for structure-based single-target drug design. DiffLinker [20] is a diffusion-based model for linker design with given fragment poses. LinkerNet [12] is a diffusion-based model for co-designing molecular fragment poses and the linker. DiffLinker and LinkerNet are repurposed for dual-target drug design as we introduced in Section 3.3.
@@
We evaluate generated ligand molecules from the perspectives of target binding affinity and molecular properties. We employ AutoDock Vina [10] to estimate the target binding affinity, following Peng et al. [37], Guan et al. [14]. We first evaluate Pocket2Mol and TargetDiff under the single-target setting as a preliminary verification (see Appendix B). For dual-target drug design, we use each method to design 10 molecules for each pair of targets, denoted as ${\mathcal{P}}_{1}$ and ${\mathcal{P}}_{2}$. (For reference molecule, Pocet2Mol and TargetDiff, the ligand molecule is generated for ${\mathcal{P}}_{1}$ but the target binding affinity is evaluated on both ${\mathcal{P}}_{1}$ and ${\mathcal{P}}_{2}$.) We then collect all generated molecules across 12,917 pairs of targets and report the mean and median (denoted as “Avg.” and “Med.” respectively) of affinity-related metrics (P-1 Vina Dock, P-2 Vina Dock, Max Vina Dock, and Dual High Affinity) and property-related metrics (drug-likeness QED [3], synthesizability SA [11], and diversity). Vina Dock incorporates a re-docking step to assess the highest binding affinity achievable. Here we introduce P-1 Vina Dock and P-2 Vina Dock to represent the Vina Dock score evaluated on ${\mathcal{P}}_{1}$ and ${\mathcal{P}}_{2}$, respectively. Besides, we introduce Max Vina Dock, which represents the maximum Vina Dock of a given molecule towards ${\mathcal{P}}_{1}$ and ${\mathcal{P}}_{2}$. The Vina Dock will be low if and only if the molecule can bind to both targets simultaneously, which is the goal of dual-target drug design. Additionally, we report Dual High Affinity (abbreviated as Dual High Aff.) which represents the proportion of generated molecules that exhibit binding affinity that exceeds that of the reference molecules on both the respective targets. This reflects the success rate in achieving higher binding affinities simultaneously on both targets in dual-target drug design. We also evaluate the RMSD between docked poses towards dual targets.
@@ 4.2 Main Results
As shown in Figure 2, DualDiff performs better than LinkerNet the RMSD between docked poses towards dual targets. This indicates that the molecules generated by DualDiff can bind to dual targets with smaller conformation change. Additionally, we provide visualization of examples of generated molecules in Figure 3. See more examples in Appendix C.
@@ Acknowledgments
This work is supported by the National Science and Technology Major Project (2023ZD0120901) and the National Natural Science Foundation of China No. 62377030.
Appendix A Term Definitions in Drug Synergy
@@
Appendix B Preliminary Verification on Single-Target Setting
@@
Appendix C Visualization of More Examples
@@
Appendix D Proof of SE(3)-Equivariance
@@
Appendix E Discussion, Limitation, and Future Work
Our work provides a novel dataset and a general framework for dual-target drug design to the community. And our method can be easily adapted to the multi-target scenario. Our work is the first step towards generative dual-target drug design. There are still limitations in our work. For example, we do not consider flexibility of proteins in our work, which is a more practical setting, though this is also a issue for most works in SBDD. We will leave it as a future work.
Appendix F Societal Impacts

### v2

@@ 1 Introduction
A promising paradigm of rational drug design is structure-based drug design (SBDD) [1], which uses computational chemistry tools in which the 3D structure of a protein target is used as the basis to identify or design new chemical entities. The foundation of structure-based drug design has been grounded in the lock-and-key hypothesis, positing that an optimal ligand molecule should possess a structure that is complementary to the target site. Recently, dual-target drug design, which aims to design “one key” for “two locks”, has attracted significant attention. Precisely, dual-target drug design [4] is a strategy in pharmaceutical research that aims to design single ligand molecules capable of interacting with two different biological targets simultaneously. A dual-target drug can potentially lower the odds of resistance developing [62] and effectively manage the disease which involve complex biological pathways with multiple proteins [49]. Recent years have witnessed a noticeable increase in the FDA’s approval of dual-target drugs [32, 34]. Please refer to Appendix A for a more comprehensive understanding of motivation, significance, and current practices of dual-target drug design.
@@ 2 Related Work
Molecular linker design, which enables the connection of molecular fragments to form potent compounds, is an effective approach in rational drug discovery. Approaches like DeLinker [26] and Develop [27] design linkers by utilizing molecular graphs with distance and angle information between anchor atoms, but they lack 3D structural information of molecules. More recent techniques, such as 3DLinker [22] and DiffLinker [25], generate linkers directly in 3D space using conditional VAEs and diffusion models, respectively, but they assume known fragment poses. LinkerNet [17] relaxes this assumption by co-designing molecular fragment poses and the linker, making it applicable in cases where fragment poses are unknown, such as in the linker design of PROTACs (PROteolysis TArgeting Chimeras). Since pharmacophore combination is a traditional strategy to design dual-target drugs, we repurpose linker design methods as strong baselines for dual-target drug design. Please refer to Appendix B for extended related works.
@@ 3.1 Data Curation
To collect drug combination pairs, we start from DrugCombDB [^1] http://drugcombdb.denglab.org/main [35]. DrugCombDB is a comprehensive database devoted to the curation of drug combinations from various data sources including high-throughput screening (HTS) assays, manual curations from the literature, FDA Orange Book and external databases. DrugCombDB comprises a total of 448,555 combinations of drugs, encompassing 2,887 unique drugs and 124 human cancer cell lines. Particularly, DrugCombDB has more than 6,000,000 quantitative dose responses, from which we determine whether a drug combination is synergistic or not. Specifically, a drug combination with positive zero interaction potency (ZIP), Bliss, Loewe and the highest single agent (HSA) scores simultaneously in at least one cell line is supposed to be a synergistic one. Please refer to Appendix C for a comprehensive understanding of these scores.
@@ 3.2 Reprogramming Target-Specific Diffusion Models for Dual-Target Drug Design
We name the compositional reverse sampling with the above SE(3)-equivariant message at each layer as DualDiff. The proof of SE(3)-equivariance can be found in Appendix G. This more meticulous composition is supposed to lead to higher-quality samples.
@@ 3.3 Repurposing Linker Design Methods for Dual-Target Drug Design
Pharmacophore combination is a prevalent strategy in traditional dual-target drug design, requiring the specialized knowledge of chemists. To automate this procedure, We design a strategy to identity crucial fragments from reference molecules of dual targets in our dataset and apply linker design methods [25, 17] to link the fragments and obtain complete molecules. Please refer to Appendix D for a more detailed justification for the choice of baselines.
@@ 4.1 Experimental Setup
We compare our method with various baselines: Pocket2Mol [45] generates 3D molecules atom by atom in an autoregressive manner given a specific protein binding site. TargetDiff [19] is a diffusion-based method which generates atom coordinates and atom types in a non-autoregressive way. Note that Pocket2Mol and TargetDiff are both proposed for structure-based single-target drug design. DiffLinker [25] is a diffusion-based model for linker design with given fragment poses. LinkerNet [17] is a diffusion-based model for co-designing molecular fragment poses and the linker. DiffLinker and LinkerNet are repurposed for dual-target drug design as we introduced in Section 3.3. The code is available at https://github.com/zhouxiangxin1998/DualDiff.
@@
We evaluate generated ligand molecules from the perspectives of target binding affinity and molecular properties. We employ AutoDock Vina [12] to estimate the target binding affinity, following Peng et al. [45], Guan et al. [19]. We first evaluate Pocket2Mol and TargetDiff under the single-target setting as a preliminary verification (see Appendix E). For dual-target drug design, we use each method to design 10 molecules for each pair of targets, denoted as ${\mathcal{P}}_{1}$ and ${\mathcal{P}}_{2}$. (For reference molecule, Pocet2Mol and TargetDiff, the ligand molecule is generated for ${\mathcal{P}}_{1}$ but the target binding affinity is evaluated on both ${\mathcal{P}}_{1}$ and ${\mathcal{P}}_{2}$.) We then collect all generated molecules across 12,917 pairs of targets and report the mean and median (denoted as “Avg.” and “Med.” respectively) of affinity-related metrics (P-1 Vina Dock, P-2 Vina Dock, Max Vina Dock, and Dual High Affinity) and property-related metrics (drug-likeness QED [3], synthesizability SA [14], and diversity). Vina Dock incorporates a re-docking step to assess the highest binding affinity achievable. Here we introduce P-1 Vina Dock and P-2 Vina Dock to represent the Vina Dock score evaluated on ${\mathcal{P}}_{1}$ and ${\mathcal{P}}_{2}$, respectively. Besides, we introduce Max Vina Dock, which represents the maximum Vina Dock of a given molecule towards ${\mathcal{P}}_{1}$ and ${\mathcal{P}}_{2}$. The Vina Dock will be low if and only if the molecule can bind to both targets simultaneously, which is the goal of dual-target drug design. Additionally, we report Dual High Affinity (abbreviated as Dual High Aff.) which represents the proportion of generated molecules that exhibit binding affinity that exceeds that of the reference molecules on both the respective targets. This reflects the success rate in achieving higher binding affinities simultaneously on both targets in dual-target drug design. We also evaluate the RMSD between docked poses towards dual targets.
@@ 4.2 Main Results
As shown in Figure 2, DualDiff performs better than LinkerNet the RMSD between docked poses towards dual targets. This indicates that the molecules generated by DualDiff can bind to dual targets with smaller conformation change. We also benchmark the inference time of baslines and our methods in Appendix H. Additionally, we provide visualization of examples of generated molecules in Figure 3. See more examples in Appendix F.
@@ Acknowledgments
This work is supported by the National Science and Technology Major Project (2023ZD0120901) and the National Natural Science Foundation of China No. 62377030. We also thank the reviewers for their valuable feedback.
Appendix A Motivation, Significance, and Current Practices of Dual-Target Drug Design
Motivation and significance of dual-target drug design
The reviewers and audience can refer to Giordano and Petrelli [16], Boran and Iyengar [5] and [41] for the concept of dual-target (or multi-target) drugs. Dual-target drugs have various advantages, especially in aspects such as anti-tumor and overcoming drug resistance.
It has been revealed that single-target drugs may not always induce the desired effect on the entire biological system even if they successfully inhibit or activate a specific target [47]. One reason is that organisms can affect effectiveness through compensatory ways. Another reason is the development of resistance either by self-modification of the target through mutation or by the adoption of new pathways by a cancer cell, for the growth and multiplication [47]. Cancer and those nervous and cardiovascular system diseases are complicated, thereby promoting the multi-targeted therapies as a better pathway to achieve the desired treatment. For example, the multiple tyrosine kinase inhibitor imatinib [^4] https://en.wikipedia.org/wiki/Imatinib induces better anti-cancer effects compared with that of gefitinib [^5] https://en.wikipedia.org/wiki/Gefitinib, which involves a single target, further indicating that drugs with multiple targets may exhibit a better chance of affecting the complex equilibrium of whole cellular networks than drugs that act on a single target. Refer to Lu et al. [41] for more examples. Notably, combination therapy also performs well in overcoming drug resistance [15]. However, under some circumstances, the combination therapy might have additive and even synergistic effects in theory; however, it often leads to some unpredictable side effects, such as increased toxicity. Thus, multi-target drugs are a better choice.
Current practices for dual-target drug design
Refer to Sun et al. [57] for a comprehensive review. The mainstream strategies are drug repurposing, skeleton modification, and linked/merged/fused pharmacophores, which involve complicated pipelines and require domain knowledge from experts. To the best of our knowledge, we are the first one to formulate dual-target drug design as a generative task and provide simple yet effective solutions.
Appendix B Extended Related Works
Our work is also relevant to neural model reprogramming [^6] Note that the reprogramming here is not relevant to the biological concept “cell reprogramming” [6].. We will discuss the connections as follows:
Elsayed et al. [13] proposed adversarial reprogramming. Specifically, given a model trained on a task which maps $x$ to $f(x)$, the adversary aims to repurpose the model for a new task which maps $g(\tilde{x})$ for inputs $\hat{x}$, by learning mapping functions $h_{f}(\cdot;\theta)$ and $h_{g}(\cdot;\theta)$ to make $h_{g}(f(h_{f}(\hat{x})))$ approximate $g(\hat{x})$. They successfully demonstrated their methods on image classification tasks. For example, a trained adversarial program can cause a classifier trained on ImageNet to an MNIST classifier.
Yang et al. [61] reprograms acoustic models for time series classification, through input transformation learning and output label mapping, which are similar to $h_{f}(\cdot)$ and $h_{g}(\cdot)$, respectively. The motivation is that data scarcity hinders researchers from using large-scale deep learning models for time-series tasks while many large-scale pre-trained speech processing models are available. Melnyk et al. [43] also follows this paradigm to reprogram pretrained language models, BERT trained on English corpus, for the antibody sequence infilling task.
The above works all aim to reprogram a model trained on a task to perform a new task. Our work also falls into this paradigm, since we focus on reprogramming models trained for single-target drug design to design dual-target drugs. And we also have a similar motivation of Yang et al. [61] and Melnyk et al. [43], because there is no training data for dual-target drug design. Differently, we focus on reprogramming a diffusion model equipped with SE(3)-equivariance which has not been widely explored in previous works due to its complex sampling process during inference. And, notably, all the above works require additional training to learn the additional parameters of input transformation (i.e., $h_{f}(\cdot;\theta)$ in Elsayed et al. [13]) for reprogramming. However, our framework works in a zero-shot manner, which means we only need to modify the sampling process of the diffusion model to perform the novel task without any additional training. Specifically, our proposed pocket alignment resembles the “input transformation”, i.e., $h_{f}(\cdot)$, and the compositional sampling resembles the “output label mapping”, i.e., $h_{g}(\cdot)$, but no learnable parameters are introduced for reprogramming. Yang et al. [61] provided theoretical results on selecting a pre-trained model for reprogramming, but it is not easy to directly apply the results to our case due to the intricate properties of diffusion models.
Appendix C Term Definitions in Drug Synergy
@@
Appendix D Justification for the Choice of Baselines
One of the traditional strategies for dual-target drug design (refer to Sun et al. [57] as a comprehensive review) is linking pharmacophores by domain experts, such as chemists or pharmaceutical scientists. We use SOTA deep-learning-based linker design methods (DiffLinker [25] and LinkerNet [19]) to mimic this procedure. Since each target pocket has a corresponding ligand molecule in our curated dataset, we can break the ligand molecule into fragments and select the ones critical for binding as pharmacophores. In this case, given respective critical fragments of dual pockets, we can design linkers and resemble them into complete molecules. Ideally, the designed molecules contain critical pharmacophores that account for binding to dual targets and serve as a potential candidate for dual-target drugs.
Appendix E Preliminary Verification on Single-Target Setting
@@
Appendix F Visualization of More Examples
@@
Appendix G Proof of SE(3)-Equivariance
@@
Appendix H Computational Efficiency
We test the time for generating 10 molecules by TargetDiff (single-target drug design) and DiffLinker, LinkerNet, CompDiff, and DualDiff (dual-target drug design). The results are shown in Table 5.
Table 5: Time Cost for generating 10 molecules by TargetDiff (single-target drug design) and DiffLinker, LinkerNet, CompDiff, and DualDiff (dual-target drug design).
Setting Method Time (seconds) Single-Target TargetDiff 220 Dual-Target DiffLinker 31 LinkerNet 30 CompDiff 395 DualDiff 493
Linker design methods (DiffLInker and LinkerNet) are faster than de novo design methods. The reasons are twofold: 1. For linker design, parts of the molecules are provided; thus, it is only necessary to generate the remaining part (i.e., the linker). 2. For DiffLinker, only one of the dual targets can serve as the input of the model. For linkerNet, it cannot use pockets as conditions. Thus, the computational cost is lower.
As the table shows, the inference time of CompDiff and DualDiff (dual-target drug design) is about twice as that of TargetDiff (single-target drug design). This is as expected because there are two heterogeneous graphs and the messages on the two graphs are processed and aggregated separately (sequentially in our implementation) and then composed as described in our paper. There is space for optimizing the inference speed of CompDiff and DualDiff by parallelling the message-passing operations over the two graphs (when GPU memory is sufficient). This code optimization will only speed up the inference and not change the generated results. Ideally, this will make the inference time of CompDiff and DualDiff almost the same as that of TargetDiff.
Though the computational cost of our method is higher than the baselines, it is still acceptable in practice.
General acceleration methods (e.g., pruning, quantization, and distillation) for deep learning models can be applied to the models used in our framework. Besides, fast diffusion sampling solvers can also be applied to our framework, with DPM-Solver [40] as one of the most representative works, which can significantly reduce the number of sampling steps without sacrificing the quality of generated samples.
Appendix I Discussion, Limitation, and Future Work
Our work provides a novel dataset and a general framework for dual-target drug design to the community. And our method can be easily adapted to the multi-target scenario. Our work is the first step towards generative dual-target drug design. There are still limitations in our work. For example, we do not consider flexibility of proteins in our work, which is a more practical setting, though this is also an issue for most works in SBDD. Another limitation of our work is the lack of validation through wet-lab experiments. We will leave these as future work.
Appendix J Societal Impacts